Solve.
(x + 7)squared = 13

Answer by stanbon(75887)

(x + 7)squared = 13
Take the square root of both sides to get:
x+7 = sqrt(13) or x+7 = -sqrt(13)
x = -7+sqrt(13) or x = -7-sqrt(13)
===================================
Cheers,
Stan H.